# Human-Like Trajectories Generation via Receding Horizon Tracking Applied to the TickTacking Interface

### v1

@@ Abstract
D. Masti, S. Menchetti, Ç. Erdem, G. Gnecco and D. Rocchesso, ”Human-Like Trajectories Generation via Receding Horizon Tracking Applied to the TickTacking Interface” in 2025 IEEE International Conference on Systems, Man and Cybernetics (SMC). IEEE, 2025.
@@
@inproceedings{masti2025human, title = {Human-Like Trajectories Generation via Receding Horizon Tracking Applied to the TickTacking Interface}, author= {Masti, Daniele and Menchetti, Stefano and Erdem, Çağrı and Gnecco, Giorgio and Rocchesso, Davide}, booktitle={2025 IEEE International Conference on Systems, Man, and Cybernetics (SMC)}, year = {2025}, organization = {IEEE} }
@@ 3 Human-Inspired Trajectory Generation Using Receding Horizon Tracking
[…]
@@ 3.1 Human-Like RH Control
This approximation is twofold: with respect to the negative log formulation of the Gaussian PDF, we neglect the additional constant term $-\log(\sqrt{2\pi}\sigma_{|\mathbf{u}|})$ and drop the squared standard deviation at the denominator of $\mu_{2}$. By combining this term with $\mu_{1}$, we ensure that velocity magnitudes remain within human-like ranges.
@@
$\left\{\begin{array}[]{ll}d_{s}\leq 0.9&\Rightarrow v_{x}=v_{x,\text{old}}\,,\\ d_{s}\geq-0.9&\Rightarrow v_{y}=v_{y,\text{old}}\,,\end{array}\right.$

### v2

@@ Abstract
D. Masti, S. Menchetti, Ç. Erdem, G. Gnecco and D. Rocchesso, ”Human-Like Trajectories Generation via Receding Horizon Tracking Applied to the TickTacking Interface” 2025 IEEE International Conference on Systems, Man, and Cybernetics (SMC), Vienna, Austria, 2025, pp. 1476-1481.
doi: 10.1109/SMC58881.2025.11342906
@@
@INPROCEEDINGS{11342906, author={Masti, Daniele and Menchetti, Stefano and Erdem, Çağrı and Gnecco, Giorgio and Rocchesso, Davide}, booktitle={2025 IEEE International Conference on Systems, Man, and Cybernetics (SMC)}, title={Human-Like Trajectories Generation via Receding Horizon Tracking Applied to the TickTacking Interface}, year={2025}, volume={}, number={}, pages={1476-1481}, doi={10.1109/SMC58881.2025.11342906} } }
@@ 3 Human-Inspired Trajectory Generation Using Receding Horizon Tracking
$\hskip 10.0pt\textrm{s.t.}\hskip 10.00002pt\mathbf{g}\left(\mathbf{S}(t),\mathbf{U}(t),\mathbf{U}_{\textrm{old}}(t),\mathbf{d}_{s}(t)\right)\leq\mathbf{0}\,$,
@@ 3.1 Human-Like RH Control
This approximation is twofold: with respect to the negative log formulation of the Gaussian PDF, we neglect the additional constant term $\log(\sqrt{2\pi}\sigma_{|\mathbf{u}|})$ and drop the squared standard deviation at the denominator of $\mu_{2}$. By combining this term with $\mu_{1}$, we ensure that velocity magnitudes remain within human-like ranges.
@@
[…]